# BiMix: Bivariate Data Mixing Law for Language Model Pretraining

### v3

@@ Abstract
Large language models have demonstrated remarkable capabilities across various tasks, primarily attributed to the utilization of diversely sourced data. However, the impact of pretraining data composition on model performance remains poorly understood. This paper introduces BiMix, a novel bivariate data mixing law that models the joint scaling behavior of domain proportions and data volume in LLM pretraining. BiMix provides a systematic framework for understanding and optimizing data mixtures across diverse domains. Through extensive experiments on two large-scale datasets, we demonstrate BiMix’s high accuracy in loss extrapolation (mean relative error $<$ 0.2%) and its generalization to unseen mixtures (R2 $>$ 0.97). Optimization of domain proportions yields superior model performance compared to existing methods. Furthermore, we establish entropy-based measures as efficient proxies for data mixing, offering a computationally lightweight strategy. Our work contributes both theoretical insights into data mixing dynamics and practical tools for enhancing LLM training efficiency, paving the way for more effective scaling strategies in language model development.
@@ 1 Introduction
The key contributions are summarized as follows:
  • A mathematically formulated mixing law that jointly models the scaling behavior of domain proportions and total training volume, providing good interpretability and functional extensibility.
  • Comprehensive experiments demonstrating the effectiveness of BiMix in predicting and optimizing model performance across diverse datasets and training scenarios.
  • Empirical evidence supporting the use of entropy-based measures as lightweight mixing proxies, offering new insights into efficient data mixture optimization.
@@ 3.1 Formulation
We propose BiMix as a system of equations to model the vectorized scaling laws of data mixing across the $m$ related domains:
@@ 5.3 Optimizing Domain Proportions for Improved Performance
To validate the efficacy of this approach, we adopt a strategy similar to that in DoReMi (Xie et al., 2023), whereby the optimized proportions obtained from smaller models are utilized to train a larger model with 1 billion parameters. After the training phase, we evaluate the model’s performance across several downstream NLP tasks. Figure 4 illustrates the variation in average model performance as the number of training steps increases. Notably, the model trained on the BiMix-optimized domain proportions consistently outperforms both the Baseline and the mixtures optimized by DoReMi. As training progresses, there remains potential for continued performance improvement. A detailed comparison of the performance on each task is included in Appendix B.
@@ Appendix A Entropy Proxies
Shannon Entropy (SE)$H_{i}(\mathcal{D}_{i})=-\sum_{x\in X_{i}}P(x)\log P(x)$, where $X_{i}$ is the set of all available tokens in domain $\mathcal{D}_{i}$ and $P(x)$ denotes the probability of observing token $x$. This proxy quantifies the expected information content associated with token appearances in the dataset, indicative of the corpus diversity.
Joint Entropy (JE)$H_{i}(X_{i}^{(t)},X_{i}^{(t+1)})=-\sum_{x\in{X}_{i}^{(t)}}\sum_{x^{\prime}\in{X}_{i}^{(t+1)}}P(x,x^{\prime})\log P(x,x^{\prime})$, where $X_{i}^{(t)}$ and $X_{i}^{(t+1)}$ represent the sets of tokens at positions $t$ and $t+1$ across all sequences in domain $\mathcal{D}_{i}$, respectively. The joint probability function $P(x,x^{\prime})$ was statistically estimated by observing a token $x$ at position $t$, followed by a token $x^{\prime}$ at position $t+1$. This metric measures the average uncertainty associated with consecutive token pairs and highlights the sequential dependencies in the dataset.
Conditional Entropy (CE)$H_{i}(X_{i}^{(t+1)}\mid X_{i}^{(t)})=-\sum_{x\in X_{i}^{(t)}}\sum_{x^{\prime}\in X_{i}^{(t+1)}}P(x,x^{\prime})\log P(x^{\prime}\mid x)$, with $X_{i}^{(t)}$, $X_{i}^{(t+1)}$, and $P(x,x^{\prime})$ as previously defined. The term $P(x^{\prime}\mid x)$ denotes the conditional probability of observing a token $x^{\prime}$ at position $t+1$ given the presence of token $x$ at position $t$. This measures the anticipated level of surprise when predicting the next token in a sequence, providing a clearer understanding of the text’s predictability and its complex linguistic structure.
Von Neumann Entropy (VNE)  In physics, the von Neumann entropy extends the concept of Gibbs entropy from classical statistical mechanics to quantum statistical mechanics. For a quantum-mechanical system described by a density matrix $\rho$, the von Neumann entropy for domain $\mathcal{D}_{i}$ is defined as
@@ Appendix C Complexity Analysis
Table 3: Complexity comparison between two mixing laws
@@
Appendix D Additional Experiments on 1B Model

### v4

@@ Abstract
Large language models have demonstrated remarkable capabilities across various tasks, primarily attributed to the utilization of diversely sourced data. However, the impact of pretraining data composition on model performance remains poorly understood. This paper introduces BiMix, a novel bivariate data mixing law that models the joint scaling behavior of domain proportions and data volume in language model pretraining. BiMix provides a systematic framework for understanding and optimizing data mixtures across diverse domains. Through extensive experiments on two large-scale datasets, we demonstrate BiMix’s high accuracy in loss extrapolation (mean relative error <0.2%) and its generalization to unseen mixtures (R2 >0.97). Optimization of domain proportions yields superior model performance compared to existing methods. Furthermore, we establish entropy-based measures as efficient proxies for data mixing, offering a computationally lightweight strategy. Our work contributes both insights into data mixing dynamics and practical tools for enhancing training efficiency, paving the way for more effective scaling strategies in language model development.
@@ 1 Introduction
The key contributions of this work are summarized as follows:
  • We introduce a novel bivariate mixing law BiMix for language model pretraining, which reveals the predictable effects of both data quantity and mixing proportions, providing significant interpretability and extensibility.
  • We offer insights into efficient mixture optimization, demonstrating that entropy measures serve as effective proxies, and provide practical guidance for data mixing.
  • We validate the utility and precision of BiMix through comprehensive experiments, highlighting its effectiveness in predicting and optimizing model performance across diverse datasets and training scenarios.
@@ 3.1 Formulation
We propose BiMix as a system of equations to model the vectorized scaling laws of data mixing across the $m$ related domains in terms of the domain proportions $\accentset{\varrightharpoonup}{r}$ and training steps $s$:
@@ 5.3 Optimizing Domain Proportions for Improved Performance
To validate the efficacy of this approach, we adopt a strategy similar to that in DoReMi (Xie et al., 2023), whereby the optimized mixtures derived from smaller models are utilized to train a larger model with billion-level parameters. Figure 4 illustrates how downstream performance varies as the number of training steps increases. Recent advancements in mixture optimization are also included for comparison, including RegMix (Liu et al., 2024) on the Pile and DoGE (Fan et al., 2023) on SlimPajama. The models trained on the BiMix-optimized mixtures demonstrated performance advantages throughout the training process. While we note that RegMix achieved performance comparable to that of BiMix, it is crucial to highlight that our approach not only optimizes mixtures but also provides a mathematical model for understanding mixing behavior. A detailed comparison of performance across each task is included in Appendix B.
@@ Appendix A Entropy Proxies
Shannon Entropy (SE)
$H_{i}(\mathcal{D}_{i})=-\sum_{x\in X_{i}}P(x)\log P(x)$, where $X_{i}$ is the set of all available tokens in domain $\mathcal{D}_{i}$ and $P(x)$ denotes the probability of observing token $x$. This proxy quantifies the expected information content associated with token appearances in the dataset, indicative of the corpus diversity.
Joint Entropy (JE)
$H_{i}(X_{i}^{(t)},X_{i}^{(t+1)})=-\sum_{x\in{X}_{i}^{(t)}}\sum_{x^{\prime}\in{X}_{i}^{(t+1)}}P(x,x^{\prime})\log P(x,x^{\prime})$, where $X_{i}^{(t)}$ and $X_{i}^{(t+1)}$ represent the sets of tokens at positions $t$ and $t+1$ across all sequences in domain $\mathcal{D}_{i}$, respectively. The joint probability function $P(x,x^{\prime})$ was statistically estimated by observing a token $x$ at position $t$, followed by a token $x^{\prime}$ at position $t+1$. This metric measures the average uncertainty associated with consecutive token pairs and highlights the sequential dependencies in the dataset.
Conditional Entropy (CE)
$H_{i}(X_{i}^{(t+1)}\mid X_{i}^{(t)})=-\sum_{x\in X_{i}^{(t)}}\sum_{x^{\prime}\in X_{i}^{(t+1)}}P(x,x^{\prime})\log P(x^{\prime}\mid x)$, with $X_{i}^{(t)}$, $X_{i}^{(t+1)}$, and $P(x,x^{\prime})$ as previously defined. The term $P(x^{\prime}\mid x)$ denotes the conditional probability of observing a token $x^{\prime}$ at position $t+1$ given the presence of token $x$ at position $t$. This measures the anticipated level of surprise when predicting the next token in a sequence, providing a clearer understanding of the text’s predictability and its complex linguistic structure.
Von Neumann Entropy (VNE)
In physics, the von Neumann entropy extends the concept of Gibbs entropy from classical statistical mechanics to quantum statistical mechanics. For a quantum-mechanical system described by a density matrix $\rho$, the von Neumann entropy for domain $\mathcal{D}_{i}$ is defined as
@@ Appendix C Complexity Analysis
Table 3: Complexity comparison of mixing laws.
@@
Appendix D Comparision with Recent Mixing Law